OpenStudy (anonymous):

If (44)x = 432, what is the value of x? 4 7 8 28

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

??? @pooja195 can you explain again

pooja195 (pooja195):

Im not sure i would say divide but y not try plugging your options in place of x?

OpenStudy (anonymous):

none of them equal 432

OpenStudy (anonymous):

@texaschic101 ????

OpenStudy (shamim):

44x/44=?

pooja195 (pooja195):

Ok then divide and round to the nearest whole number

pooja195 (pooja195):

So i was right before.....

OpenStudy (anonymous):

im sorry im dumb but what do i have to do?

OpenStudy (anonymous):

I think its mostly likely 8

OpenStudy (shamim):

44x=432 x=432/44 x=?

OpenStudy (shamim):

U know 44x/44=?

OpenStudy (shamim):

Tell me 2/2=?

OpenStudy (anonymous):

omg im so sorry its if \[ 4^{4} \] to the power of x = \[4^{32}\] what is the value of x?

OpenStudy (shamim):

Tell me first 44/44=?

OpenStudy (anonymous):

@pooja195 @shamim

OpenStudy (shamim):

I think u did not understand ur previous question. Right????

OpenStudy (anonymous):

this is the same question but i wrote it wrong

OpenStudy (anonymous):

@shamim are you going to help me?

OpenStudy (anonymous):

@pooja195 @shamim @texaschic101 please help me!!!!